### v1

# SARCLIP: A Vision Language Foundation Model for Semantic Understanding and Target Recognition in SAR Imagery

Abstract
Synthetic Aperture Radar (SAR) has emerged as a crucial imaging modality due to its all-weather capabilities. While recent advancements in self-supervised learning and Masked Image Modeling (MIM) have paved the way for SAR foundation models, these approaches primarily focus on low-level visual features, often overlooking multimodal alignment and zero-shot target recognition within SAR imagery. To address this limitation, we construct SARCLIP-1M, a large-scale vision language dataset comprising over one million text-image pairs aggregated from existing datasets. We further introduce SARCLIP, the first vision language foundation model tailored for the SAR domain. Our SARCLIP model is trained using a contrastive vision language learning approach by domain transferring strategy, enabling it to bridge the gap between SAR imagery and textual descriptions. Extensive experiments on image-text retrieval and zero-shot classification tasks demonstrate the superior performance of SARCLIP in feature extraction and interpretation, significantly outperforming state-of-the-art foundation models and advancing the semantic understanding of SAR imagery. The code and datasets will be released soon.
I Introduction
SAR is a high-resolution imaging technology with all-weather, day-night operability and strong penetration capabilities, widely applied in military, environmental, maritime, and disaster monitoring tasks. Compared with optical imagery, SAR images are characterized by speckle noise, geometric distortions, and limited semantic textures, which pose significant challenges for downstream tasks such as ship detection [1], aircraft recognition [2], object detection [3], semantic segmentation [4], and image classification [5]. These challenges highlight the need for robust and generalizable feature representations specifically tailored to SAR data.
Figure 1: Comparison of image-text datasets in remote sensing. Gray circles represent datasets in the optical image domain, blue circles denote those in the SAR domain, and the red circle indicates our proposed SARCLIP-1M. In the SAR domain, SARCLIP-1M significantly surpasses existing datasets in both image and text volume.
Recent developments in vision foundation models (VFMs) have led to promising generalization capabilities across domains. As illustrated in Fig. 2, existing VFMs can be broadly categorized into three paradigms: (a) contrastive learning (CL-based) methods such as SimCLR [6], which aim to learn discriminative representations by aligning augmented views of the same image; (b) masked image modeling (MIM-based) methods that reconstruct occluded image regions to learn spatial representations; and (c) CLIP-based approaches [7] that align images and texts via contrastive loss on large-scale paired data, showing remarkable transferability in cross-modal tasks.
Figure 2: The paradigm for foundation model: (a) CL-based methods, (b) MIM-based methods, (c) CLIP-based methods.
The self-supervised training paradigm has recently driven the development of several vision foundation models in remote sensing, including RS-BYOL [8], ScaleMAE [9], and Cross-Scale MAE [10]. Notably, ScaleMAE [9] focuses on learning multi-scale representations by reconstructing both low- and high-frequency components across defined spatial scales. Similarly, Cross-Scale MAE [10] enforces consistency across scales using a combination of contrastive and generative losses to enhance self-supervised remote sensing representations. However, these methods predominantly address the extraction of visual features. Inspired by CLIP [7], RemoteCLIP [11] pioneered the first vision language foundation model (VLFM) in remote sensing, aiming to capture semantic features by aligning image-text pairs. Subsequent works like RS5M [12] and SkyScript [13] have further explored this aspect. However, these models are pre-trained on optical image, leaving the SAR modality underexplored due to its distinct characteristics and lack of large-scale image-text datasets in SAR domain.
In recent years, the growing interest in SAR modality has led to the creation of various large-scale SAR datasets [14, 15, 16, 17]. Specifically, SARDet-100K consolidates existing SAR datasets to form a large collection encompassing ships, aircraft, cars, bridges, tanks, and harbors. Building upon this, SARATR-X extends SARDet-100K by integrating additional classification datasets to create the larger SARDet-180K. Both SARDet-100K [14] and SARATR-X [15] then employ MIM training strategy to establish their respective foundation models. Furthermore, recent advanced methods have to address the application of multi-modal LLMs in the SAR field. Particularly, SARChat-2M [18] introduces a substantial benchmark with two million multimodal dialogues, facilitating intelligent interpretation of SAR imagery through LLM-based conversational paradigms. A comparable initiative has been undertaken with SARLANG-1M [19]. However, despite their impressive performance, existing SAR vision foundation models and multimodal LLMs remain limited by the lack of textual annotations in current datasets, hindering their ability to fully capture rich semantic information.
To bridge this gap, we propose SARCLIP, the first CLIP-based vision-language foundation model designed specifically for SAR imagery. As shown in Fig. 1, we construct SARCLIP-1M, a large-scale SAR image-text dataset comprising 1.7 million pairs across diverse object categories and land cover types. These pairs are generated by leveraging domain knowledge, spatial rules, and templated text synthesis strategies. Based on SARCLIP-1M dataset, we adopt a two-stage domain strategy to transfer knowledge from optical to SAR domain, enabling semantic alignment between modalities. As shown in Fig. 3, this approach enables the extraction of comprehensive general features from SAR images and significantly enhances their semantic understanding, thereby elevating performance for SAR interpretation.
Figure 3: (a) Examples from the SARCLIP-1M dataset; (b) Domain transfer training strategy for SARCLIP.
The key contributions of our SARCLIP are summarized as follows:
  • We propose SARCLIP-1M, a novel large-scale vision-language dataset containing 1.7 million image-text pairs, encompassing various object and land cover types.
  • To the best of our knowledge, SARCLIP is the first VLFM tailored for SAR imagery, enabling SAR-specific cross-modal alignment, target recognition and zero-shot classification.
  • Extensive experiments demonstrate that SARCLIP consistently surpasses existing state-of-the-art VLFMs on multiple downstream tasks, highlighting its strong generalization and semantic representation capabilities in the SAR domain.
II Related Work
II-A Vision Language Model for Remote Sensing
Vision-language models have significantly advanced remote sensing through the development of multi-modal techniques. These methods broadly fall into two categories: contrastive methods and generative methods. Among contrastive approaches, CLIP [7] stands out as a pioneering work, employing a two-tower architecture to align visual and language features through extensive data from Internet. Inspired by CLIP, several studies have adapted this paradigm to remote sensing, yielding models such as RemoteCLIP [11], GeoRSCLIP [12], SkyScript [13], RSMCLIP [20] and Mall et al. [21]. Notably, RemoteCLIP [11] enhances pre-training by converting detection and segmentation annotations into image captions, facilitating CLIP-style contrastive learning for image-text alignment in remote sensing. In the realm of generative methods, works like GeoChat [22], EarthGPT [23], and LHRS-Bot [24] implement auto-regressive large language model (LLM) architectures for vision-text alignment. For instance, GeoChat [22] introduces a multimodal model with LLaVa [25] architecture in remote sensing, achieving multi-granularity alignment through parameter-efficient fine-tuning. However, most of these existing methods primarily focus on the optical modality in remote sensing, largely overlooking investigations in the SAR domain.
II-B SAR Foundation Model
Foundation models, by pretraining on large-scale data, are capable of capturing generalizable visual representations that effectively support a wide array of downstream tasks. In remote sensing, recent advancements have seen the utilization of self-supervised learning and MIM techniques, as demonstrated by approaches such as RingMo [26], SatMAE [27], ScaleMAE [9], OmniSat [28], and Cross-Scale MAE [10]. These models have been widely applied in tasks like aerial object detection and target recognition.
With the increasing availability of SAR imagery, a multitude of datasets have emerged, including SARDet-100K [14], SARATR-X [15], SAR-JEPA [29], FAIR-CSAR [16], and ATRNet-STAR [17]. Specifically, SARDet-100K constructs SAR foundation model through MIM training, initially pretraining on aerial view images before transferring to SAR imagery. Building upon this, SARATR-X leverages various classification datasets to build the SARDet-180K dataset, subsequently yielding a MIM-based SAR foundation model. Furthermore, SARLANG-1M [19] utilizes LLMs for SAR image interpretation, while SARChat [18] supports key tasks such as visual understanding and object detection in SAR imagery. Nevertheless, a common limitation among these existing SAR foundation models is their primary focus on low-level image features, often failing to capture deeper semantic information and realize multi-modal alignment within SAR images.
TABLE I: Illustration of SARCLIP-1M, which includes 8 SAR datasets. Cls:Classification. Det.:Detection. Cap.:Caption. # Train Imgs.: Number of training imges. # Train Caps.: Number of training captions. # Val Imgs.: Number of validation images. # Val Caps.: Number of validation captions. # Test Pairs: Number of image-text pairs in testing set.
Datasets Year Task # Train Imgs. # Train Caps. # Val Imgs. # Val Caps. # Test Pairs.
MSTAR [30] 1995 Cls. 3,046 15,230 9,855 49,275 180
SARSim [31] 2017 Cls. 21,168 105,840
OpenSARShip [32] 2017 Cls. 26,679 133,395
SAMPLE [33] 2019 Cls. 5,380 26,900
ATRNet-STAR [17] 2025 Cls.&Det. 68,091 340,455 29,284 146,420 6,667
SARDet-100K [14] 2024 Det. 94,493 472,465 10,492 52,460 2,783
FAIR-CSAR [16] 2024 Det. 51,948 259,740 11,790 58,950 7,096
SARLANG-1M-Captions [19] 2025 Cap. 9,191 31,968 3,939 13,682 3,902
SARCLIP-1M(ours) 2025 Cap. 279,996 1,385,993 65,360 320,787 20,628
TABLE II: Overview of the datasets comprising SARCLIP-1M dataset. The three datasets1 are SpaceNet6 [34], DFC2023 [35], and OpenEarthMap [36]
. Datasets Descriptions MSTAR Contains X-band SAR imagery of military vehicles. SARSim A simulation dataset providing vehicle samples across 7 categories. OpenSARShip Features ship slices derived from European C-band Sentinel-1 satellite data. SAMPLE A public X-band SAR dataset of 10 vehicle classes, with synthetic and real image pairs. ATRNet-STAR A large-scale SAR dataset offering 40 fine-grained vehicle target classes. SARDet-100K Compiled from 10 existing SAR detection datasets, encompassing 5 object classes. FAIR-CSAR A large-scale, fine-grained SLC SAR dataset covering 22 subcategories. SARLANG-1M-Captions Contains over 45,000 SAR image captions based on three datasets1. SARCLIP-1M (ours) Comprises over 1.7 million image-text pairs, including ship, vehicle, aircraft, and other land covers.
III Methodology
This section introduce the paradigm of our framework, dataset construction approach and training strategy for SARCLIP.
III-A Problem Definition
In this section, we investigate the paradigm of learning joint representations from SAR images and their corresponding textual descriptions. Specifically, we construct SARCLIP-1M dataset $\mathcal{D}=\{(\mathbf{I}_{i},\mathbf{T}_{i})\}_{i=1}^{M}$ consisting of SAR images $\mathbf{I}_{i}\in\mathcal{R}^{H\times W}$ with the corresponding descriptions $\mathbf{T}_{i}\in\mathcal{T}$. As shown in Fig. 3, our objective is to learn a pair of modality-specific encoders that project both SAR images and text into a shared semantic space. Specifically, we define a visual encoder $f_{v}:\mathcal{R}^{H\times W}\rightarrow\mathcal{R}^{d}$ that maps the input SAR image to a $d$-dimensional visual feature embedding $z_{v}^{i}=f_{v}({\mathbf{I}_{i}})$, and a textual encoder $f_{t}:\mathcal{R}^{\mathcal{T}}\rightarrow\mathcal{R}^{d}$ that maps the textual input to a corresponding textual embedding $z_{t}^{i}=f_{t}({\mathbf{T}_{i}})$. The goal is to align the embeddings $z_{v}^{i}$ and $z_{t}^{i}$ of matched image-text pairs in a common representation space, such that semantically similar inputs across modalities are embedded close to each other. This formulation enables the model to bridge the modality gap between SAR images and natural language, thereby supporting a wide range of downstream tasks such as cross-modal retrieval, target recognition, and zero-shot classification in SAR domains.
III-B SARCLIP-1M Dataset Construction
To address the aforementioned challenges, we build a large-scale vision-language dataset SARCLIP-1M collected from existing classification, detection and caption datasets in SAR domain. As illustrated in Table I and II, several large datasets have released in the past two year. For classification datasets, we employ 10 simple and complex description template to generate image captions. As for detection, we additionally design template to describe object absolute position and relative spatial position between targets in the images. The details for template design method are as follows:
(1) General descriptions. We utilize simple templates such as ”A SAR image of the [class]” where [class] is replaced with a category name from classification datasets, or object type and quantity from detection datasets.
(2) Complex descriptions. We employ complex templates like ”A SAR image reveals the distinct texture and structure of the [class].” where the usage of template is the same as general. These complex templates can enhance the diversity of descriptions and the robustness of models
(3) Absolute region descriptions. The image is divided into five regions: upper left, upper right, bottom left, bottom right and center region. We calculate the IoU between the box annotations of target and each region to determine its location. Templates like ”A SAR image of [classes] located in the [location] of the image.” are utilized.
Figure 4: The word cloud of SARCLIP-1M dataset.
(4) Relative region descriptions. Templates are utilized to describe spatial relationships between targets. Relative templates like ”In this SAR image, the [class1] in the [location1] are positioned [relative_direction] the [class2] in the [location2].” In these template, ”[relative_direction]” includes above, below, left and right.
Based on the above templates design, we obtain caption for each SAR imagery. Then, A large language model is utilized to verify the fluency and grammatical correctness of text. Finally, we build our SARCLIP-1M dataset contains 279,996 images and 1,385,993 captions in training set, and 20,628 image-text pairs in testing set. As shown in Fig. 4, the dataset includes diverse target types such as ship, vehicle, aircraft, bridge and other land cover types.
III-C Model Training
III-C1 SARCLIP training.
To further enhance the semantic understanding ability of the model for SAR image, we utilize two stages training strategy to train our SARCLIP. Firstly, we train vision large model on natural remote sensing dataset like RemoteCLIP which contains LoveDA [37], DOTA [38], RSCID [39] and others. By this strategy, our model is capable of better understanding and recognizing the unique characteristics of remote sensing imagery, particularly in identifying small scale and densely distributed objects. Secondly, we finetune this pretrained model on our SARCLIP-1M dataset to transfer knowledge from natural to SAR domain.
In both two stage, we train our model using a contrastive learning framework that encourages positive pairs $(z_{v}^{i},z_{t}^{i})$ to be more similar than negative pairs. The objective is to minimize the following symmetric InfoNCE-based contrastive loss which can be define as follows:
$\mathcal{L}_{\text{clip}}=-\frac{1}{N}\sum_{i=1}^{N}\biggl\{\bigg[$ $\log\frac{\exp(\text{sim}(z_{v}^{i},z_{t}^{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(z_{v}^{i},z_{t}^{j})/\tau)}$
$+$ $\log\frac{\exp(\text{sim}(z_{t}^{i},z_{v}^{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(z_{t}^{i},z_{v}^{j})/\tau)}\bigg]/2\biggr\}$ (1)
where $\text{sim}(a,b)=\frac{a^{\top}b}{\|a\|\|b\|}$ is cosine similarity, $\tau$ donates a temperature hyperparameter, and $N$ is the batch size.
III-C2 Downstream fine-tuning.
To adapt our pretrained SARCLIP model to the SAR target recognition task, a lightweight fine-tuning is designed. Specifically, we freeze the parameters of the visual encoder and train a task-specific classification head $f_{h}:\mathcal{R}^{d}\rightarrow\mathcal{R}^{C}$ on the top of frozen visual features, the formula are as follows:
$z_{v}=f_{v}(I)$ (2)
$\hat{y}=f_{h}(z_{v})$ (3)
where $I$ is the input SAR image, $C$ denotes the number of target categories, and $z_{v}\in\mathcal{R}^{C}$ is the visual feature from SARCLIP. The model is optimized using the standard cross-entropy loss:
$\mathcal{L}_{\mathrm{cls}}=-\log\frac{\exp(\hat{y}_{y})}{\sum_{c=1}^{C}\exp(\hat{y}_{c})}$ (4)
where $y$ is the ground-truth label of the input image.
TABLE III: Retrieval performance on the SARCLIP-1M test set and improvements over the SARCLIP baseline(%). LAION is an optical image dataset in the natural scene domain, while RS5M, Skyscript, and R3+D10+S4 are optical image datasets in the remote sensing domain. , §, and denote models pretrained on RS5M, Skyscript, and R3+D10+S4, respectively.
Image to Text Text to Image Mean
Method Image Backbone Pretrain Data Tune On R@1 R@5 R@10 R@1 R@5 R@10 Recall
OpenCLIP ResNet-50 LAION 0.01 0.02 0.07 0.01 0.04 0.13 0.04
OpenCLIP ViT-B-32 LAION 0.02 0.05 0.10 0.02 0.12 0.20 0.08
OpenCLIP ViT-L-14 LAION 0.01 0.07 0.13 0.05 0.16 0.30 0.12
GeoRSCLIP ViT-B-32 LAION RS5M 0.03 0.09 0.18 0.06 0.19 0.31 0.14
GeoRSCLIP ViT-L-14 LAION RS5M 0.05 0.15 0.24 0.09 0.32 0.51 0.22
GeoRSCLIP ViT-H-14 LAION RS5M 0.01 0.06 0.12 0.08 0.32 0.51 0.18
SkyCLIP ViT-B-32 LAION SkyScript 0.01 0.08 0.15 0.05 0.23 0.37 0.14
SkyCLIP ViT-L-14 LAION SkyScript 0.03 0.10 0.20 0.14 0.39 0.65 0.25
RemoteCLIP ResNet-50 LAION R3+D10+S4 0.01 0.07 0.16 0.03 0.12 0.21 0.10
RemoteCLIP ViT-B-32 LAION R3+D10+S4 0.02 0.07 0.13 0.03 0.14 0.24 0.10
RemoteCLIP ViT-L-14 LAION R3+D10+S4 0.01 0.08 0.16 0.05 0.17 0.33 0.13
SARCLIP ResNet-50 LAION SARCLIP-1M 2.71 10.82 17.57 3.07 11.76 19.03 10.49
SARCLIP ResNet-50 R3+D10+S4 SARCLIP-1M 2.98 10.79 17.35 3.21 12.02 19.39 10.95 (+4.38)
HarMA ViT-B-32 LAION SARCLIP-1M 1.15 4.80 9.00 1.13 5.05 9.21 5.06
SARCLIP ViT-B-32 LAION SARCLIP-1M 3.42 12.82 20.32 3.63 13.72 21.40 12.55
SARCLIP ViT-B-32 RS5M SARCLIP-1M 3.73 13.33 20.61 3.95 13.99 21.77 12.89
SARCLIP § ViT-B-32 SkyScript SARCLIP-1M 3.79 13.23 20.31 3.72 13.80 21.46 12.72
SARCLIP ViT-B-32 R3+D10+S4 SARCLIP-1M 3.58 13.25 20.89 3.79 13.64 21.52 12.77 (+1.75)
SARCLIP ViT-L-14 LAION SARCLIP-1M 4.73 15.53 23.81 4.75 16.09 24.51 14.90
SARCLIP ViT-L-14 RS5M SARCLIP-1M 4.81 15.60 23.57 4.77 16.56 25.26 15.09
SARCLIP § ViT-L-14 SkyScript SARCLIP-1M 4.69 15.70 23.84 4.67 16.34 24.65 14.98
SARCLIP ViT-L-14 R3+D10+S4 SARCLIP-1M 4.69 15.87 24.04 4.90 16.62 25.42 15.25 (+2.34)
IV Experiment
IV-A Datasets
To evaluate the superiority of our proposed SARCLIP, three extensively adopted SAR target recognition datasets and our SARCLIP-1M test retrieval dataset is utilized for downstream performance evaluation.
IV-A1 SARCLIP-1M test.
This SAR retrieval dataset comprises 20,628 image-text pairs compiled from existing datasets. The test set consists of 180 images from MSTAR, 6,667 images from ATRNet-STAR, 2,783 images from SARDet-100K, 7,096 images from FAIR-CSAR, and 3,902 images from SARLANG-1M-Captions. Each image is uniquely captioned.
IV-A2 MSTAR-SOC.
This SAR target recognition dataset is acquired by an X-band radar operating in HH polarization mode with a resolution of 0.3 meters. It comprises 10 military vehicle targets and is partitioned into 4 experimental settings, as established by [40]. Specifically, The SOC setting has 2,747 training images (17 angles) and 2,425 test images (15 angles), with all 10 classes present in both sets.
IV-A3 FUSAR-ship.
It is a dataset designed for ship recognition, comprising 15 main ship classes, 98 subclasses, and various non-ship maritime targets [15]. It is built from 126 Gaofen-3 images captured in ultrafine resolution mode (1.124 $\times$ 1.728 m) with dual-polarization (DH and DV). The dataset covers a wide range of scenarios, including open sea, coastal areas, rivers, islands, and land backgrounds.
IV-A4 SAR-VSA.
This SAR target recognition dataset comprises 25 fine-grained categories, integrating data from MSTAR, FUSAR-ship , and SAR-ACD [41]. It contains 11,045 training images and 8,161 testing images, as constructed by [15].
IV-B Evaluation Metrics
We employ image text retrieval and target recognition as our downstream task to demonstrate the effectiveness of our framework. For the retrieval task, we adopt Recall at top-$K$ ($\mathrm{R@K}$), where $K\in\{1,5,10\}$. The formula can be defined as:
$\mathrm{R@K}=\frac{1}{N}\sum_{i=1}^{N}I\left[y_{i}\in\mathrm{TopK}(q_{i})\right]$ (5)
where $N$ is the number of queries, $y_{i}$ is the ground-truth item for query $q_{i}$, and ${I}[\cdot]$ is the indicator function. For the target recognition task, we utilize accuracy ($\mathrm{ACC}$) as metric which can be defined as:
$\mathrm{ACC}=\frac{1}{N}\sum_{i=1}^{N}{I}\left[\hat{y}_{i}=y_{i}\right]$ (6)
where $\hat{y}_{i}$ is the predicted result and $y_{i}$ is the ground-truth label.
TABLE IV: Recognition results on MSTAR-SOC and SAR-VSA dataset and improvements over the OpenCLIP(%)($\mathbf{\%}$).
dataset
Method Backbone Param. MSTAR-SOC SAR-VSA
OpenCLIP ResNet-50 38 70.76 71.25
RemoteCLIP ResNet-50 38 66.60 69.83
SARCLIP ResNet-50 38 99.75 78.75
SARCLIP ResNet-50 38 99.54(+40.67) 81.08(+13.79)
OpenCLIP ViT-B-32 87 61.89 71.60
GeoRSCLIP ViT-B-32 87 76.00 74.85
RemoteCLIP ViT-B-32 87 69.97 73.56
SARCLIP ViT-B-32 87 99.54 87.30
SARCLIP ViT-B-32 87 99.58 86.93
SARCLIP § ViT-B-32 87 99.62 87.56
SARCLIP ViT-B-32 87 99.58(+60.89) 87.77(+22.58)
OpenCLIP ViT-L-14 304 78.80 81.58
GeoRSCLIP ViT-L-14 304 82.26 80.07
GeoRSCLIP ViT-H-14 632 79.84 78.39
RemoteCLIP ViT-L-14 304 78.80 80.02
SARCLIP ViT-L-14 304 99.54 88.92
SARCLIP ViT-L-14 304 99.62 89.75
SARCLIP § ViT-L-14 304 99.71 88.56
SARCLIP ViT-L-14 304 99.83(+26.68) 88.83(+8.88)
IV-C Implementation Details
We develop SARCLIP based on the OpenCLIP framework. Automatic mixed-precision (AMP) training is employed to reduce memory usage. We adopt ResNet-50, ViT-B-32, and ViT-L-14 as image backbones, with learning rates set to 5e-4, 5e-5, and 5e-5, respectively. ResNet-50 is trained for 30 epochs, while ViT-B-32 and ViT-L-14 are trained for 10 epochs. The batch size is set to 256. Training is accelerated using the Adam optimizer, combined with a linear warm-up and cosine learning rate schedule. Downstream experiments are conducted using the wise-ft [42] framework. For the downstream target recognition task, we freeze the image backbone and fine-tune only the linear classification layers for 10,000 epochs. All experiments are conducted on two 80GB NVIDIA H100 GPUs.
IV-D Comparing with other Methods
In this section, we compare out method with other state-of-the-art CLP-based method include OpenCLIP, RemoteCLIP, GeoRSCLIP, SkyCLIP and HarMA [43].
TABLE V: Zero-shot classification results on FUSAR-ship and improvements over the RemoteCLIP baseline $(\mathbf{\%})$.
Method Backbone Param. FUSAR-ship
SARCLIP ResNet-50 38 15.51
SARCLIP ResNet-50 38 14.69
RemoteCLIP ViT-B-32 87 1.19
SARCLIP ViT-B-32 87 8.97
SARCLIP ViT-B-32 87 11.25
SARCLIP § ViT-B-32 87 3.12
SARCLIP ViT-B-32 87 0.51
RemoteCLIP ViT-L-14 304 9.72
SARCLIP ViT-L-14 304 6.32
SARCLIP ViT-L-14 304 8.16
SARCLIP § ViT-L-14 304 11.12
SARCLIP ViT-L-14 304 12.41(+27.67)
IV-D1 Retrieval results on SARCLIP-1M test dataset.
As presented in Table III, we compare our SARCLIP method with several state-of-the-art VLFMs designed for remote sensing. The results clearly indicate that existing models exhibit poor performance in SAR image-text retrieval tasks. In contrast to standard CLIP models (e.g., OpenCLIP), our training strategy effectively transfers knowledge from the optical domain (e.g. RS5M) to the SAR domain (SARCLIP-1M). Specifically, SARCLIP achieves a 15.25% mean recall while SARCLIP at 14.90%, demonstrating a significant improvement and highlighting the efficacy of our approach in multi-modal alignment in SAR domain.
TABLE VI: Ablation study result on several downstream task. Img. Enc. and Text Enc. donate the image encoder and text decoder respectively.(%)($\mathbf{\%}$)
Model Architecture SARCLIP-1M test SOC VSA FUSAR
Img. Enc. Text Enc. Mean Recall ACC ACC ACC
5.23 50.72 78.48 1.49
9.57 99.62 88.94 10.85
15.25 99.83 88.83 12.41
Figure 5: Feature space visualization of SARCLIP image encoder on three downstream datasets (ViT-L-14).
IV-D2 Target recognition results on MSTAR-SOC and SAR-VSA dataset.
Table IV presents the performance of our SARCLIP models on downstream target recognition tasks after fine-tuning a linear layer, thereby demonstrating their visual understanding capabilities. The experimental results show that our SARCLIP series models consistently achieve superior performance. On the MSTAR-SOC dataset, SARCLIP achieves an impressive 99.83% accuracy, significantly outperforming optical VLFMs which typically hover around 80%. Similarly, on the SAR-VSA dataset, our model reaches 89.75% accuracy (specifically, SARCLIP with ViT-L-14), notably surpassing OpenCLIP’s 81.58%. These results underscore the strong visual feature extraction and understanding capabilities of our SARCLIP models for SAR imagery.
IV-D3 Zero-shot classification results on FUSAR-ship dataset.
Table V presents zero-shot classification results on the FUSAR-ship dataset, highlighting the generalization ability of SARCLIP without fine-tuning. While RemoteCLIP performs poorly (e.g., 1.19% for ViT-B-32), our SARCLIP models exhibit notably stronger zero-shot capabilities. In particular, SARCLIP with a ViT-L-14 backbone achieves the best performance at 12.41%, demonstrating improved transferability to unseen SAR ship categories. These results suggest that pre-training on a diverse SAR vision-language dataset like SARCLIP-1M significantly enhances zero-shot generalization over models trained primarily on visible-light data.
Figure 6: Ablation study on training layers of SARCLIP on the SARCLIP-1M test set (Mean Recall %).
IV-E Ablation Studies
IV-E1 Effect of the training encoder.
This section presents an ablation study evaluating the contributions of the image and text encoders to model performance across various downstream tasks. As summarized in Table VI, when only the text encoder is used, the model performs poorly, achieving just 5.23% mean recall on the SARCLIP-1M test set and 1.49% accuracy on FUSAR, highlighting the limited discriminative power of text alone for SAR imagery understanding. In contrast, activating only the image encoder significantly improves performance, with mean recall reaching 9.57%, and accuracy on MSTAR-SOC and SAR-VSA increasing to 99.62% and 88.94%, respectively demonstrating the image encoder’s effectiveness in SAR feature extracting. The combination of both encoders yields the best results, achieving 15.25% mean recall on SARCLIP-1M and 99.83% accuracy on SOC. Notably, zero-shot accuracy on FUSAR improves to 12.41%, confirming that cross-modal contrastive learning is essential for aligning semantic representations and enhancing generalization in both retrieval and zero-shot classification tasks.
IV-E2 Effect of the training layer.
As shown in Fig. 6, we evaluate the impact of varying the number of activated layers in the image and text encoders on the SARCLIP-1M test set. The results reveal a consistent trend: model performance improves as more layers are activated. This highlights the importance of deep, expressive representations in both modalities for effective SAR image-text understanding.
IV-F Visualization
As shown in Fig. 5, we utilize t-SNE [44] technique to visualize features extracted by SARCLIP (ViT-L-14) on three downstream datasets. The visualizations plots exhibit well-separated clusters across classes, demonstrating the model’s strong feature discrimination and providing insights into its visual and semantic representations.
V Conclusion
We present SARCLIP-1M, a large-scale SAR image-text dataset, and propose SARCLIP, the first vision-language foundation model for SAR. Trained via a two-stage domain transfer strategy, SARCLIP effectively transfers knowledge from optical remote sensing data to SAR imagery, achieving strong multi-modal alignment, target recognition and zero-shot capability. In the future, we will explore integrating multi-modal large language models and agent techniques for SAR image interpretation.