### v2

# SARVLM: A Vision Language Foundation Model for Semantic Understanding and Target Recognition in SAR Imagery

Abstract
Synthetic Aperture Radar (SAR) is a crucial imaging modality thanks to its all-weather capability. Although recent advances in self-supervised learning and masked image modeling (MIM) have enabled SAR foundation models, these methods largely emphasize low-level visual features and often overlook multimodal alignment and zero-shot target recognition in SAR imagery. To address this, we construct SARVLM-1M, a large-scale vision–language dataset with over one million image–text pairs aggregated from existing datasets. We further propose a domain transfer training strategy to mitigate the large gap between natural and SAR imagery. Building on this, we develop SARVLM, the first vision language foundation model (VLM) tailored to SAR, comprising SARCLIP and SARCap. SARVLM is trained with a vision–language contrastive objective under the proposed domain transfer strategy, bridging SAR imagery and textual descriptions. Extensive experiments on image text retrieval, zero-shot classification, semantic localization, and imagery captioning demonstrate that SARVLM delivers superior feature extraction and interpretation, outperforming state-of-the-art VLMs and advancing SAR semantic understanding. Code and datasets will be released soon.
1 Introduction
SAR is a high-resolution imaging technology with all-weather, day-night operability and strong penetration capabilities, widely applied in military, environmental, maritime, and disaster monitoring tasks. Compared with optical imagery, SAR images are characterized by speckle noise, geometric distortions, and limited semantic textures, which pose significant challenges for downstream tasks such as ship detection [SSDD], aircraft recognition [SFRNet], object detection [RSAR], semantic segmentation [liu2025learning], and image classification [deng2024hyperspectral]. These challenges highlight the need for robust and generalizable feature representations specifically tailored to SAR data.
Figure 1: Comparison of image-text datasets in remote sensing. Gray circles represent datasets in the optical image domain, blue circles denote those in the SAR domain, and the red circle indicates our proposed SARVLM-1M. In the SAR domain, SARVLM-1M significantly surpasses existing datasets in both image and text volume.
Recent developments in vision foundation models (VFMs) have led to promising generalization capabilities across domains. As illustrated in Figure 2, existing VFMs can be broadly categorized into three paradigms: (a) contrastive learning (CL-based) methods such as SimCLR [SimCLR], which aim to learn discriminative representations by aligning augmented views of the same image; (b) masked image modeling (MIM-based) methods that reconstruct occluded image regions to learn spatial representations; and (c) CLIP-based approaches [CLIP] that align images and texts via contrastive loss on large-scale paired data, showing remarkable transferability in cross-modal tasks.
Figure 2: The paradigm for foundation model: (a) CL-based methods, (b) MIM-based methods, (c) CLIP-based methods.
The self-supervised training paradigm has recently driven the development of several vision foundation models in remote sensing, including RS-BYOL [RS-BYOL], ScaleMAE [ScaleMAE], and Cross-scale MAE [Cross-Sclae-MAE]. Notably, ScaleMAE [ScaleMAE] focuses on learning multi-scale representations by reconstructing both low- and high-frequency components across defined spatial scales. Similarly, Cross-scale MAE [Cross-Sclae-MAE] enforces consistency across scales using a combination of contrastive and generative losses to enhance self-supervised remote sensing representations. However, these methods predominantly address the extraction of visual features. Inspired by CLIP [CLIP], RemoteCLIP [RemoteCLIP] pioneered the first vision language foundation model (VLFM) in remote sensing, aiming to capture semantic features by aligning image-text pairs. Subsequent works like RS5M [RS5M] and SkyScript [Skyscript] have further explored this aspect. However, these models are pre-trained on optical image, leaving the SAR modality underexplored due to its distinct characteristics and lack of large-scale image-text datasets in SAR domain.
Figure 3: (a) Examples from the SARVLM-1M dataset; (b) Two-stage domain transfer training strategy for SARCLIP.
In recent years, the growing interest in SAR modality has led to the creation of various large-scale SAR datasets [SARDet-100K, SARATR, FAIR-CSAR, ATRNet-STAR]. Specifically, SARDet-100K consolidates existing SAR datasets to form a large collection encompassing ships, aircraft, cars, bridges, tanks, and harbors. Building upon this, SARATR-X extends SARDet-100K by integrating additional classification datasets to create the larger SARDet-180K. Both SARDet-100K [SARDet-100K] and SARATR-X [SARATR] then employ MIM training strategy to establish their respective foundation models. Furthermore, recent advanced methods have to address the application of multi-modal LLMs in the SAR field. Particularly, SARChat-2M [SARChat-Bench-2M] introduces a substantial benchmark with two million multimodal dialogues, facilitating intelligent interpretation of SAR imagery through LLM-based conversational paradigms. A comparable initiative has been undertaken with SARLANG-1M [SARLANG-1M]. However, despite their impressive performance, existing SAR vision foundation models and multimodal LLMs remain limited by the lack of textual annotations in current datasets, hindering their ability to capture rich semantic information fully.
To bridge this gap, we propose SARVLM consisting of SARCLIP and SARCap, the first vision-language foundation model designed specifically for SAR imagery. As shown in Figure 1, we construct SARVLM-1M, a large-scale SAR image-text dataset comprising 1.7 million pairs across diverse object categories and land cover types. These pairs are generated by leveraging domain knowledge, spatial rules, and templated text synthesis strategies. Based on SARVLM-1M dataset, we adopt a two-stage domain strategy to transfer knowledge from optical to SAR domain, enabling semantic alignment between modalities. As shown in Figure 3, this approach enables the extraction of comprehensive general features from SAR images and significantly enhances their semantic understanding, thereby elevating performance for SAR interpretation.
The key contributions of our SARVLM are summarized as follows:
  • We propose SARVLM-1M, a novel large-scale vision-language dataset containing 1.7 million image-text pairs, encompassing various object and land cover types.
  • To the best of our knowledge, SARVLM is the first VLFM tailored for SAR imagery, enabling SAR-specific cross-modal alignment, target recognition, zero-shot classification, semantic localization, and imagery captioning.
  • Extensive experiments demonstrate that SARVLM consistently surpasses existing state-of-the-art VLFMs on multiple downstream tasks, highlighting its strong generalization and semantic representation capabilities in the SAR domain.
2 Related Work
2.1 Vision Language Model for Remote Sensing
Vision-language models have significantly advanced remote sensing through the development of multi-modal techniques. These methods broadly fall into two categories: contrastive methods and generative methods. Among contrastive approaches, CLIP [CLIP] stands out as a pioneering work, employing a two-tower architecture to align visual and language features through extensive data from Internet. Inspired by CLIP, several studies have adapted this paradigm to remote sensing, yielding models such as RemoteCLIP [RemoteCLIP], GeoRSCLIP [RS5M], SkyScript [Skyscript], RSMCLIP [rethinking] and Mall et al. [mall2023remote]. Notably, RemoteCLIP [RemoteCLIP] enhances pre-training by converting detection and segmentation annotations into image captions, facilitating CLIP-style contrastive learning for image-text alignment in remote sensing. In the realm of generative methods, works like GeoChat [Geochat], EarthGPT [EarthGPT], and LHRS-Bot [LHRS-Bot] implement auto-regressive large language model (LLM) architectures for vision-text alignment. For instance, GeoChat [Geochat] introduces a multimodal model with LLaVa [LLaVa] architecture in remote sensing, achieving multi-granularity alignment through parameter-efficient fine-tuning. However, most of these existing methods primarily focus on the optical modality in remote sensing, largely overlooking investigations in the SAR domain.
2.2 SAR Foundation Model
Foundation models, by pretraining on large-scale data, are capable of capturing generalizable visual representations that effectively support a wide array of downstream tasks. In remote sensing, recent advancements have seen the utilization of self-supervised learning and MIM techniques, as demonstrated by approaches such as RingMo [RingMo], SatMAE [SatMAE], ScaleMAE [ScaleMAE], OmniSat [OmniSat], and Cross-scale MAE [Cross-Sclae-MAE]. These models have been widely applied in tasks like aerial object detection and target recognition.
With the increasing availability of SAR imagery, a multitude of datasets have emerged, including SARDet-100K [SARDet-100K], SARATR-X [SARATR], SAR-JEPA [SAR-JEPA], FAIR-CSAR [FAIR-CSAR], and ATRNet-STAR [ATRNet-STAR]. Specifically, SARDet-100K constructs SAR foundation model through MIM training, initially pretraining on aerial view images before transferring to SAR imagery. Building upon this, SARATR-X leverages various classification datasets to build the SARDet-180K dataset, subsequently yielding a MIM-based SAR foundation model. Furthermore, SARLANG-1M [SARLANG-1M] utilizes LLMs for SAR image interpretation, while SARChat [SARChat-Bench-2M] supports key tasks such as visual understanding and object detection in SAR imagery. Nevertheless, a common limitation among these existing SAR foundation models is their primary focus on low-level image features, often failing to capture deeper semantic information and realize multi-modal alignment within SAR images.
Table 1: Illustration of SARVLM-1M, which includes 8 SAR datasets. Cls:Classification. Det.:Detection. Cap.:Caption. # Train Imgs.: Number of training imges. # Train Caps.: Number of training captions. # Val Imgs.: Number of validation images. # Val Caps.: Number of validation captions. # Test Pairs: Number of image-text pairs in testing set.
Datasets Year Task # Train Imgs. # Train Caps. # Val Imgs. # Val Caps. # Test Pairs.
MSTAR [MSTAR] 1995 Cls. 3,046 15,230 9,855 49,275 180
SARSim [SARSim_1, SARSim_2] 2017 Cls. 21,168 105,840
OpenSARShip [OpenSARship] 2017 Cls. 26,679 133,395
SAMPLE [SAMPLE] 2019 Cls. 5,380 26,900
ATRNet-STAR [ATRNet-STAR] 2025 Cls.&Det. 68,091 340,455 29,284 146,420 6,667
SARDet-100K [SARDet-100K] 2024 Det. 94,493 472,465 10,492 52,460 2,783
FAIR-CSAR [FAIR-CSAR] 2024 Det. 51,948 259,740 11,790 58,950 7,096
SARLANG-1M-Captions [SARLANG-1M] 2025 Cap. 9,191 31,968 3,939 13,682 3,902
SARVLM-1M(ours) 2025 Cap. 279,996 1,385,993 65,360 320,787 20,628
Figure 4: The framework of SARCap method.
3 Methodology
This section introduce the paradigm of our framework, dataset construction approach and training strategy for SARCLIP and SARCap.
3.1 Problem Definition
In this section, we investigate the paradigm of learning joint representations from SAR images and their corresponding textual descriptions. Specifically, we construct SARVLM-1M dataset $\mathcal{D}=\{(\mathbf{I}_{i},\mathbf{T}_{i})\}_{i=1}^{M}$ consisting of SAR images $\mathbf{I}_{i}\in\mathcal{R}^{H\times W}$ with the corresponding descriptions $\mathbf{T}_{i}\in\mathcal{T}$. As shown in Figure 3, our objective is to learn a pair of modality-specific encoders that project both SAR images and text into a shared semantic space. Specifically, we define a visual encoder $f_{v}:\mathcal{R}^{H\times W}\rightarrow\mathcal{R}^{d}$ that maps the input SAR image to a $d$-dimensional visual feature embedding $z_{v}^{i}=f_{v}({\mathbf{I}_{i}})$, and a textual encoder $f_{t}:\mathcal{R}^{\mathcal{T}}\rightarrow\mathcal{R}^{d}$ that maps the textual input to a corresponding textual embedding $z_{t}^{i}=f_{t}({\mathbf{T}_{i}})$. The goal is to align the embeddings $z_{v}^{i}$ and $z_{t}^{i}$ of matched image-text pairs in a common representation space, such that semantically similar inputs across modalities are embedded close to each other. This formulation enables the model to bridge the modality gap between SAR images and natural language, thereby supporting a wide range of downstream tasks such as cross-modal retrieval, target recognition, and zero-shot classification in SAR domains.
Figure 5: The word cloud of SARVLM-1M dataset.
3.2 SARVLM-1M Dataset Construction
To address the aforementioned challenges, we build a large-scale vision-language dataset SARVLM-1M collected from existing classification, detection and caption datasets in SAR domain. As illustrated in Table 1, several large datasets have released in the past two year. For classification datasets, we employ 10 simple and complex description template to generate image captions. As for detection, we additionally design template to describe object absolute position and relative spatial position between targets in the images. The details for template design method are as follows:
(1) General descriptions. We utilize simple templates such as ”A SAR image of the [class]” where [class] is replaced with a category name from classification datasets, or object type and quantity from detection datasets.
(2) Complex descriptions. We employ complex templates like ”A SAR image reveals the distinct texture and structure of the [class].” where the usage of template is the same as general. These complex templates can enhance the diversity of descriptions and the robustness of models
(3) Absolute region descriptions. The image is divided into five regions: upper left, upper right, bottom left, bottom right and center region. We calculate the IoU between the box annotations of target and each region to determine its location. Templates like ”A SAR image of [classes] located in the [location] of the image.” are utilized.
(4) Relative region descriptions. Templates are utilized to describe spatial relationships between targets. Relative templates like ”In this SAR image, the [class1] in the [location1] are positioned [relative_direction] the [class2] in the [location2].” In these template, ”[relative_direction]” includes above, below, left and right.
Based on the above templates design, we obtain caption for each SAR imagery. Then, A large language model is utilized to verify the fluency and grammatical correctness of text. Finally, we build our SARVLM-1M dataset contains 279,996 images and 1,385,993 captions in training set, and 20,628 image-text pairs in testing set. As shown in Figure 5, the dataset includes diverse target types such as ship, vehicle, aircraft, bridge and other land cover types.
3.3 Model Training
3.3.1 SARCLIP training
To further enhance the semantic understanding ability of the model for SAR image, we utilize two stages training strategy to train our SARCLIP. Firstly, we train vision large model on natural remote sensing dataset like RemoteCLIP which contains LoveDA [LoveDA], DOTA [DOTA], RSCID [RSCID] and others. By this strategy, our model is capable of better understanding and recognizing the unique characteristics of remote sensing imagery, particularly in identifying small scale and densely distributed objects. Secondly, we finetune this pretrained model on our SARVLM-1M dataset to transfer knowledge from natural to SAR domain.
In both two stage, we train our model using a contrastive learning framework that encourages positive pairs $(z_{v}^{i},z_{t}^{i})$ to be more similar than negative pairs. The objective is to minimize the following symmetric InfoNCE-based contrastive loss which can be define as follows:
$\mathcal{L}_{\text{clip}}=-\frac{1}{N}\sum_{i=1}^{N}\biggl\{\bigg[$ $\log\frac{\exp(\text{sim}(z_{v}^{i},z_{t}^{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(z_{v}^{i},z_{t}^{j})/\tau)}$
$+$ $\log\frac{\exp(\text{sim}(z_{t}^{i},z_{v}^{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(z_{t}^{i},z_{v}^{j})/\tau)}\bigg]/2\biggr\}$ (1)
where $\text{sim}(a,b)=\frac{a^{\top}b}{\|a\|\|b\|}$ is cosine similarity, $\tau$ denotes a temperature hyperparameter, and $N$ is the batch size.
3.3.2 SARCap training
For the SAR image captioning task, we use the CoCa [Coca] architecture, which combines contrastive loss and captioning loss. The captioning loss minimizes the difference between the predicted and ground-truth captions:
$\mathcal{L}_{\text{caption}}=-\sum_{t=1}^{T}\log P(w_{t}|w_{1},...,w_{t-1},z_{v})$ (2)
where $T$ is the caption length, $w_{t}$ is the $t$-th token, and $P(\cdot)$ denotes the predicted probability of $w_{t}$ conditioned on the previous tokens and the visual feature.
3.3.3 Downstream fine-tuning
To adapt our pretrained SARCLIP model to the SAR target recognition task, a lightweight fine-tuning is designed. Specifically, we freeze the parameters of the visual encoder and train a task-specific classification head $f_{h}:\mathcal{R}^{d}\rightarrow\mathcal{R}^{C}$ on the top of frozen visual features, the formula are as follows:
$z_{v}=f_{v}(I)$ (3)
$\hat{y}=f_{h}(z_{v})$ (4)
where $I$ is the input SAR image, $C$ denotes the number of target categories, and $z_{v}\in\mathcal{R}^{C}$ is the visual feature from SARCLIP. The model is optimized using the standard cross-entropy loss:
$\mathcal{L}_{\mathrm{cls}}=-\log\frac{\exp(\hat{y}_{y})}{\sum_{c=1}^{C}\exp(\hat{y}_{c})}$ (5)
where $y$ is the ground-truth label of the input image.
Table 2: Retrieval performance on the SARVLM-1M test set and improvements over the SARCLIP baseline(%). LAION [OpenCLIP] is an optical image dataset in the natural scene domain, while RS5M, Skyscript, and R3+D10+S4 are optical image datasets in the remote sensing domain. , §, and denote models pretrained on RS5M [RS5M], Skyscript [Skyscript], and R3+D10+S4 [RemoteCLIP], respectively.
Image to Text Text to Image Mean
Method Image Backbone Pretrain Data Tune On R@1 R@5 R@10 R@1 R@5 R@10 Recall
OpenCLIP ResNet-50 LAION 0.01 0.02 0.07 0.01 0.04 0.13 0.04
OpenCLIP ViT-B-32 LAION 0.02 0.05 0.10 0.02 0.12 0.20 0.08
OpenCLIP ViT-L-14 LAION 0.01 0.07 0.13 0.05 0.16 0.30 0.12
GeoRSCLIP ViT-B-32 LAION RS5M 0.03 0.09 0.18 0.06 0.19 0.31 0.14
GeoRSCLIP ViT-L-14 LAION RS5M 0.05 0.15 0.24 0.09 0.32 0.51 0.22
GeoRSCLIP ViT-H-14 LAION RS5M 0.01 0.06 0.12 0.08 0.32 0.51 0.18
SkyCLIP ViT-B-32 LAION SkyScript 0.01 0.08 0.15 0.05 0.23 0.37 0.14
SkyCLIP ViT-L-14 LAION SkyScript 0.03 0.10 0.20 0.14 0.39 0.65 0.25
RemoteCLIP ResNet-50 LAION R3+D10+S4 0.01 0.07 0.16 0.03 0.12 0.21 0.10
RemoteCLIP ViT-B-32 LAION R3+D10+S4 0.02 0.07 0.13 0.03 0.14 0.24 0.10
RemoteCLIP ViT-L-14 LAION R3+D10+S4 0.01 0.08 0.16 0.05 0.17 0.33 0.13
SARCLIP ResNet-50 LAION SARVLM-1M 2.71 10.82 17.57 3.07 11.76 19.03 10.49
SARCLIP ResNet-50 R3+D10+S4 SARVLM-1M 2.98 10.79 17.35 3.21 12.02 19.39 10.95 (+4.38)
HarMA ViT-B-32 LAION SARVLM-1M 1.15 4.80 9.00 1.13 5.05 9.21 5.06
SARCLIP ViT-B-32 LAION SARVLM-1M 3.42 12.82 20.32 3.63 13.72 21.40 12.55
SARCLIP ViT-B-32 RS5M SARVLM-1M 3.73 13.33 20.61 3.95 13.99 21.77 12.89
SARCLIP § ViT-B-32 SkyScript SARVLM-1M 3.79 13.23 20.31 3.72 13.80 21.46 12.72
SARCLIP ViT-B-32 R3+D10+S4 SARVLM-1M 3.58 13.25 20.89 3.79 13.64 21.52 12.77 (+1.75)
SARCLIP ViT-L-14 LAION SARVLM-1M 4.73 15.53 23.81 4.75 16.09 24.51 14.90
SARCLIP ViT-L-14 RS5M SARVLM-1M 4.81 15.60 23.57 4.77 16.56 25.26 15.09
SARCLIP § ViT-L-14 SkyScript SARVLM-1M 4.69 15.70 23.84 4.67 16.34 24.65 14.98
SARCLIP ViT-L-14 R3+D10+S4 SARVLM-1M 4.69 15.87 24.04 4.90 16.62 25.42 15.25 (+2.34)
4 Experiment
4.1 Datasets
To evaluate the superiority of our proposed SARCLIP, we utilize downstream dataset as follows: SARVLM-1M test. This SAR retrieval dataset comprises 20,628 image-text pairs compiled from existing datasets. The test set consists of 180 images from MSTAR, 6,667 images from ATRNet-STAR, 2,783 images from SARDet-100K, 7,096 images from FAIR-CSAR, and 3,902 images from SARLANG-1M-Captions. Each image is uniquely captioned.
MSTAR-SOC. This SAR target recognition benchmark acquired by an X-band HH radar at 0.3 m resolution, containing 10 military vehicle classes and four standard experimental settings [AConvNet]. In the SOC setting, all 10 classes appear in 2,747 training images from 17 aspect angles and 2,425 test images from 15 angles.
FUSAR-Ship. This ship recognition dataset comprising 15 main ship classes, 98 subclasses, and diverse non-ship maritime targets [SARATR]. It is built from 126 Gaofen-3 ultrafine-resolution images (1.124 $\times$ 1.728 m, dual-pol DH/DV), covering open sea, coastal waters, rivers, islands, and land backgrounds.
SAR-VSA. This SAR target recognition dataset comprises 25 fine-grained categories, integrating data from MSTAR [MSTAR], FUSAR-ship [FUSAR-ship], and SAR-ACD [SAR-ACD]. It contains 11,045 training images and 8,161 testing images, as constructed by [SARATR].
AIR-SLT. This is a remote sensing semantic localization benchmark with 22 large-scale images and 59 text-query test cases, each paired with ground-truth bounding boxes [SeLo, SeLov2].
4.2 Evaluation Metrics
We employ image text retrieval and target recognition as our downstream task to demonstrate the effectiveness of our framework. For the retrieval task, we adopt Recall at top-$K$ ($\mathrm{R@K}$), where $K\in\{1,5,10\}$. The formula can be defined as:
$\mathrm{R@K}=\frac{1}{N}\sum_{i=1}^{N}I\left[y_{i}\in\mathrm{TopK}(q_{i})\right]$ (6)
where $N$ is the number of queries, $y_{i}$ is the ground-truth item for query $q_{i}$, and ${I}[\cdot]$ is the indicator function. For the target recognition task, we utilize accuracy ($\mathrm{ACC}$) as metric which can be defined as:
$\mathrm{ACC}=\frac{1}{N}\sum_{i=1}^{N}{I}\left[\hat{y}_{i}=y_{i}\right]$ (7)
where $\hat{y}_{i}$ is the predicted result and $y_{i}$ is the ground-truth label. For the semantic localization task, we adopt Rsu, Rda, Ras, and Rmi as evaluation metrics, following [SeLo]. For the image captioning task, we employ BLEU [bleu], ROUGE_L [rouge_L], METEOR [meteor], and CIDEr [cider].
4.3 Implementation Details
We develop SARCLIP based on the OpenCLIP framework. Automatic mixed-precision (AMP) training is employed to reduce memory usage. We adopt ResNet-50, ViT-B-32, and ViT-L-14 as image backbones, with learning rates set to 5e-4, 5e-5, and 5e-5, respectively. ResNet-50 is trained for 30 epochs, while ViT-B-32 and ViT-L-14 are trained for 10 epochs. The batch size is set to 256. Training is accelerated using the Adam optimizer, combined with a linear warm-up and cosine learning rate schedule. Downstream experiments are conducted using the wise-ft [wise-ft] and SLM[SeLo] framework. For the downstream target recognition task, we freeze the image backbone and fine-tune only the linear classification layers for 10,000 epochs. All experiments are conducted on two 80GB NVIDIA H100 GPUs. For SAR imagery caption task, we employ optical image caption dataset HQRS-210K [HQRS-210K] as pretrained data.
4.4 Comparing with other Methods
In this section, we compare our method with other state-of-the-art CLIP-based and CoCa-based method on downstream tasks.
Table 3: Recognition results on MSTAR-SOC and SAR-VSA dataset and improvements over the OpenCLIP(%)($\mathbf{\%}$).
dataset
Method Backbone Param. MSTAR-SOC SAR-VSA
OpenCLIP ResNet-50 38 70.76 71.25
RemoteCLIP ResNet-50 38 66.60 69.83
SARCLIP ResNet-50 38 99.75 78.75
SARCLIP ResNet-50 38 99.54(+40.67) 81.08(+13.79)
OpenCLIP ViT-B-32 87 61.89 71.60
GeoRSCLIP ViT-B-32 87 76.00 74.85
RemoteCLIP ViT-B-32 87 69.97 73.56
SARCLIP ViT-B-32 87 99.54 87.30
SARCLIP ViT-B-32 87 99.58 86.93
SARCLIP § ViT-B-32 87 99.62 87.56
SARCLIP ViT-B-32 87 99.58(+60.89) 87.77(+22.58)
OpenCLIP ViT-L-14 304 78.80 81.58
GeoRSCLIP ViT-L-14 304 82.26 80.07
GeoRSCLIP ViT-H-14 632 79.84 78.39
RemoteCLIP ViT-L-14 304 78.80 80.02
SARCLIP ViT-L-14 304 99.54 88.92
SARCLIP ViT-L-14 304 99.62 89.75
SARCLIP § ViT-L-14 304 99.71 88.56
SARCLIP ViT-L-14 304 99.83(+26.68) 88.83(+8.88)
Table 4: Zero-shot classification results on FUSAR-ship and improvements over the RemoteCLIP baseline $(\mathbf{\%})$.
Method Backbone Param. FUSAR-ship
SARCLIP ResNet-50 38 15.51
SARCLIP ResNet-50 38 14.69
RemoteCLIP ViT-B-32 87 1.19
SARCLIP ViT-B-32 87 8.97
SARCLIP ViT-B-32 87 11.25
SARCLIP § ViT-B-32 87 3.12
SARCLIP ViT-B-32 87 0.51
RemoteCLIP ViT-L-14 304 9.72
SARCLIP ViT-L-14 304 6.32
SARCLIP ViT-L-14 304 8.16
SARCLIP § ViT-L-14 304 11.12
SARCLIP ViT-L-14 304 12.41(+27.67)
Table 5: Semantic Localization results on AIR-SLT datase.
[…]
RemoteCLIP ViT-L-14 0.7707 0.6722 0.2876 0.7257
SARCLIP ViT-L-14 0.5643 0.4830 0.5631 0.5000
SARCLIP ViT-L-14 0.5654 0.4531 0.5601 0.4889
SARCLIP § ViT-L-14 0.5675 0.4792 0.5677 0.4981
SARCLIP ViT-L-14 0.5741 0.4173 0.5731 0.4834
4.4.1 Retrieval results on SARVLM-1M test dataset
As presented in Table 2, we compare our SARCLIP methods with several state-of-the-art VLFMs designed for remote sensing. The results clearly indicate that existing models exhibit poor performance in SAR image-text retrieval tasks. In contrast to standard CLIP models (e.g., OpenCLIP), our training strategy effectively transfers knowledge from the optical domain (e.g. RS5M) to the SAR domain (SARVLM-1M). Specifically, SARCLIP achieves a 15.25% mean recall while SARCLIP at 14.90%, demonstrating a significant improvement and highlighting the efficacy of our approach in multi-modal alignment in SAR domain.
4.4.2 Target recognition results on MSTAR-SOC and SAR-VSA dataset
Table 3 presents the performance of our SARCLIP models on downstream target recognition tasks after fine-tuning a linear layer, thereby demonstrating their visual understanding capabilities. The experimental results show that our SARCLIP series models consistently achieve superior performance. On the MSTAR-SOC dataset, SARCLIP achieves an impressive 99.83% accuracy, significantly outperforming optical VLFMs which typically hover around 80%. Similarly, on the SAR-VSA dataset, our SARCLIP reaches 89.75% accuracy, notably surpassing OpenCLIP’s 81.58%. These results underscore the strong visual feature extraction our SARCLIP models for SAR imagery.
Figure 6: Feature space visualization of SARCLIP image encoder on three downstream datasets (ViT-L-14).
Table 6: Image Caption results on SARVLM-1M test datset with image encoder(ViT-L-14)(%)($\mathbf{\%}$).
Method Image Encoder Pretrain Data Tune On BLUE-1 BLUE-2 BLUE-3 BLUE-4 METEOR ROUGE_L CIDEr
CoCa-Vanilla ViT-L-14 LAION 8.75 3.21 1.14 0.32 10.54 10.67 4.17
CoCa-HQRS ViT-L-14 LAION HQRS 10.79 4.02 1.14 0.39 8.21 13.11 7.47
SARCap ViT-L-14 SARVLM-1M 22.42 14.44 10.45 7.79 19.23 21.76 24.32
SARCap ViT-L-14 LAION+HQRS SARVLM-1M 26.10 17.13 11.93 8.47 22.18 26.15 29.58
4.4.3 Zero-shot classification on FUSAR-ship dataset
Table 4 summarizes zero-shot classification on FUSAR-ship, evaluating SARCLIP without fine-tuning. RemoteCLIP performs poorly (e.g., 1.19% for ViT-B-32), whereas SARCLIP variants show much stronger zero-shot recognition. In particular, SARCLIP with a ViT-L-14 backbone reaches 12.41%, indicating better transfer to unseen SAR ship categories. This demonstrates that pre-training on the large-scale SARVLM-1M SAR vision-language corpus greatly enhances zero-shot generalization over models trained mainly on visible-light data.
4.4.4 Semantic localization on AIR-SLT dataset
Table 5 reports semantic localization results on the optical AIR-SLT dataset under a cross-domain setting. Our SARCLIP achieves 0.5000 Rmi compared to 0.7257 for RemoteCLIP, showing that SARCLIP retains a certain level of generalization to optical scenes. As illustrated in Figure 7, given query captions such as “four white planes” and ”white airport”, our model successfully captures the plane features but misses the ”white airport” region, which is mainly attributed to the intrinsic differences between SAR and optical imagery.
4.4.5 SAR imagery caption on SARVLM-1M
Table 8 presents the caption result from our SARCap framework. After high-quality optical remote sensing fine-tuning, CoCa-HQRS performs better than CoCa-Vanilla. We utilize a progressive training strategy to obtain our SAR imagery captioning model SARCap which outperforms SARCap from scripts. As shown in Figure 8, our SARCap generate fluent and meaningful caption for different SAR sences.
4.5 Ablation Studies
4.5.1 Effect of downmain transferring strategy
There is a large perspective gap between natural and optical remote sensing images, but a smaller gap between optical remote sensing and SAR scenes, making the transfer to SAR more challenging. Validation on tasks like image-text retrieval shows the effectiveness of our progressive knowledge transfer strategy.
4.5.2 Effect of the Training Encoder
This ablation study examines the impact of image and text encoders on downstream tasks. Table 7 shows that the text encoder alone achieves poor performance, with 5.23% mean recall on SARVLM-1M and 1.49% accuracy on FUSAR. The image encoder alone improves performance to 9.57% mean recall and 99.62% accuracy on MSTAR-SOC. Combining both encoders yields the best results, with 15.25% mean recall on SARVLM-1M and 99.83% accuracy on SOC, while zero-shot accuracy on FUSAR rises to 12.41%. This confirms that cross modal contrastive learning is key for better generalization.
Figure 7: Visualization of results on semantic localization task from SARCLIP (ViT-L-14).
Figure 8: Visualization of SAR imagery caption results from SARCap.
4.5.3 Effect of the training layer
As shown in Figure 9, we evaluate the impact of varying the number of activated layers in the image and text encoders on the SARVLM-1M test set. The results reveal a consistent trend: model performance improves as more layers are activated. This highlights the importance of deep, expressive representations in both modalities for effective SAR image-text understanding.
4.6 Visualization
As shown in Figure 6, we use t-SNE [t-SNE] to visualize SARCLIP (ViT-L-14) features on three downstream datasets. The well-separated clusters indicate strong feature discrimination and meaningful semantic representations.
Figure 9: Ablation study on training layers of SARCLIP on the SARVLM-1M test set (Mean Recall %).
Table 7: Ablation study result on several downstream tasks. Img. Enc. and Text Enc. denote the image encoder and text decoder respectively.(%)($\mathbf{\%}$)
Model Architecture SARVLM-1M test SOC VSA FUSAR
Img. Enc. Text Enc. Mean Recall ACC ACC ACC
5.23 50.72 78.48 1.49
9.57 99.62 88.94 10.85
15.25 99.83 88.83 12.41
5 Conclusion
We present SARVLM-1M, a large-scale SAR image-text dataset, and propose SARCLVLM, the first vision-language foundation model for SAR. Trained via a two-stage domain transfer strategy, SARCLIP effectively transfers knowledge from optical remote sensing data to SAR imagery, achieving strong multi-modal alignment, target recognition and zero-shot and semantic localization capability while SARCap performs well in SAR imagery captioning. In the future, we will explore integrating multi-modal large language models and agent techniques for SAR image interpretation.